# Assessing the viability of non-light water reactor concepts for electricity and heat generation in decarbonized energy systems

### v1

@@ Abstract
Recent pledges to triple global nuclear capacity by 2050 suggest a "nuclear renaissance," bolstered by reactor concepts such as sodium-cooled fast reactors, high-temperature reactors, and molten salt reactors. These technologies claim to address the challenges of today’s high-capacity light-water reactors, i.e., cost overruns, delays, and social acceptance, while also offering additional non-electrical applications. However, this analysis reveals that none of these concepts currently meet the prerequisites of affordability, competitiveness, or commercial availability. We omit social acceptability. The cost analysis reveals optimistic FOAK cost assumptions of 5,623 to 9,511 USD per kW, and NOAK cost projections as low as 1,476 USD per kW. At FOAK cost, the applied energy system model includes no nuclear power capacity, and thus indicates that significant cost reductions would be required for these technologies to contribute to energy system decarbonization. In low-cost scenarios, reactors capable of producing high temperature heat become competitive with other low-carbon technologies. We conclude that, for reactor capacties to increase significantly, a focus on certain technology lines ist necessary. However, until a concept becomes viable and commercially available, policymakers should prioritize existing technologies to decarbonize energy systems.
@@
Analysis of historical and ongoing non-electrical applications of nuclear power
Assessment of the viability of different reactor concepts for such applications
Detailed cost analysis for proposed non-light water reactor concepts
Energy system modeling of nuclear cogeneration with parallel renewables expansion
@@ 1 Introduction
There is a broad consensus that decarbonizing energy systems is necessary to limit global warming and reduce the potentially hazardous effects of climate change [1]. However, extensive debates exist on the means to this end, especially regarding the technologies that are part of a cost-efficient and executable solution. In these debates, nuclear power is often considered a low-carbon technology that will play a significant part in global energy system decarbonization [2, 1], albeit existing uncertainties regarding the cost-efficiency and availability of this technology [3, 4, 5, 6, 7, 8]. In the last decade, these prospects have launched a narrative of a so-called "nuclear renaissance" coinciding with the promotion of so-called new reactor (SNR) concepts, or "advanced" reactor technologies [9, 10]. This is supported by recent governmental proclamations of plans to triple the installed nuclear power plant capacity by 2050 [11].
@@
These ambitious expansion targets stand against a number of challenges that the industry is currently facing [7], including cost escalations when building new high-capacity light water reactors (LWR) [3, 4], uncertainty regarding the availability of so-called "small modular reactors" (SMR) and other "advanced" reactor concepts [5] as well as a limited actor base [12]. These challenges raise concerns about whether the nuclear industry will be able to live up to the high expectations raised by proponents of the proclaimed nuclear renaissance.
@@
Currently, several SNR concepts are being developed. According to a 2023 report by the U.S. "Committee on Laying the Foundations for New and Advanced Nuclear Reactors" [13, p.43], these concepts must fulfill four conditions to be competitive in an energy system with a high share of fluctuating renewables: They must be 1) "affordable for owner investment without ’betting the company’", 2) "economically competitive with other technologies […]", 3) "socially acceptable", and 4) "commercially available." This coincides with required overnight construction costs in the range of 2000 to 6000 USD per kW and the establishment and diffusion of non-electrical uses for nuclear power plants [14, 15, 13]. However, there is a considerable discrepancy between cost assumptions or projections and actual costs of high capacity LWR projects, especially in OECD countries [8]. For SMR and SNR concepts, limited operational experience hinders the availability of accurate cost and other performance data [7]. Especially for SMR concepts, manufacturers’ cost projections appear somewhat optimistic compared to production theory [5]. Nonetheless, smaller sizes, non-electrical uses, and other benefits of nuclear, such as reliability and flexibility, could increase the competitiveness of both SNR and SMR concepts [16, 17, 18].
@@
Further, most previous studies limit their assessment of the role of nuclear power in future energy systems to existing technologies (i.e., high capacity LWRs) and electricity production [14, 19, 20, 8]. At the same time, some literature refers to "advanced" reactor technologies with links to (thermal) storage options (e.g., [2, 21]). However, to our knowledge, a gap exists regarding the detailed assessment of the potential of non-electrical applications for nuclear power. Further, a model-based consideration of the potential non-electrical applications of different nuclear reactor concepts is missing due to the limitation of previous studies on a single respective reactor design.
@@
Thus, in this work, we assess the potential validity of SMR and SMR concept applicability in future energy systems, including non-electrical applications, by following a three-step approach. First, based on an extensive literature assessment, we investigate potential non-electrical applications of nuclear power, e.g., for heat applications, and discuss technology availability. Second, we discuss the potential costs for each technology based on an analysis of 51 references. Third, based on the prior steps, we investigate the economic efficiency of these technologies by integrating them into a European energy system model [22] with the ability to generate both electricity and heat. Assuming SNR technologies become commercially available, we find that costs must be reduced substantially, i.e., well below current LWR project costs in Europe, and reactor concepts compete mainly for heat provision in which competitive low-carbon technologies are similarly expensive.
@@
The remainder of this work is structured as follows. Section LABEL:background provides the main results of the literature assessment on use cases and technology availability, and of the cost data analysis. Section 3 introduces the applied energy system model and relevant assumptions and data. The results of the energy system model are presented in Section 4, and are discussed in Section 5. Section 6 concludes. Further background information is provided in the Appendix.
2 Background
LABEL:background
@@ 2 Background
This section gives an overview of potential non-electrical use cases for nuclear power and of reactor concept availability. Further, we provide an assessment of cost projections for these currently commercially unavailable reactor concepts. This section provides relevant background information for the implementation of the energy system model regarding scenario definition and technology selection.
@@ 2.1 Potential Non-electrical Applications of Nuclear Power
Despite the dominance of LWRs in today’s global nuclear fleet, which is mainly used to generate electricity [23], the cogeneration of electricity and heat has been the subject of literature and technology development for decades, as in general, nuclear reactors are already "heat generation devices" [24, p. 23]. Csik and Kupitz [24] further differentiate between reactors intended for heat generation only and those that can also produce electricity, allowing some degree of flexibility. However, cogeneration experience remains limited to individual applications; see Appendix A. This results from the low-level output temperature from LWRs, which limits the economic retrievability of energy. However, this output temperature could theoretically be raised at the expense of electricity generation [25].
@@
Figure 1: Potential industrial applications for existing and future reactor types, own depiction based on [27, p. 12]
@@ 2.1.1 District Heat
Nuclear power reactors have been supplying low-temperature heat up to 100°C for district heating for decades [30, 31]. While there are issues related to acceptance, e.g., addressing the proximity to populated areas [13], regulation, e.g., regarding the establishment of emergency planning zones [31, 18], as well as technology, e.g., the challenge of efficiency losses when transferring heat over long distances [25, 17], SMRs and SNRs are thought to be especially applicable to this use case [31]. There have been several attempts to design reactors specifically for district heating in Sweden, Finland, and China [30, 31]. Mostly in Eastern Europe, there exists decade-long experience with district heating via nuclear cogeneration, for example, at the Kozloduy plant in Bulgaria, the Bohunice V-2 plant in Slovakia, or the Paks plant in Hungary. The heat is used for small demand centers nearby [32, 29]. Additionally, in 2019, two Westinghouse AP1000 reactors began providing district heat to the city of Haiyang in China [33].
@@ 2.1.2 Process Heat for Industry
Decarbonizing the provision of heat for industrial processes such as cement production and steel manufacturing is challenging as most processes require heightened temperature levels of several hundred °C and a near-constant provision of energy [34]. For decades, it has been argued that nuclear power can provide both the required heat and capacity. For example, current high-capacity LWRs can theoretically produce heat ranging from 200 to 300 °C, applicable for pulp, paper, and textiles manufacturing or seawater desalination (see Section 2.1.4), and advanced gas-cooled reactors currently operating in the U.K., can provide heat up to 650°C, a temperature level relevant for the chemical industry [24]. So far, however, experience is limited to individual, mostly low-capacity demonstration projects that use mainly excess heat from electricity generation. Examples include the Calder Hall gas-cooled reactors, built in the 1950s in the U.K., which provided steam to the nearby Windscale reprocessing plant, or the Gösgen plant in Switzerland, which has been supplying heat to a nearby cardboard factory using 1% of its steam since 1979 [35]. In Canada, the Bruce heavy-water nuclear plant provided heat for its own usage but also to a greenhouse, an ethanol plant, and a plastic film production plant, amongst others, covering distances of over six kilometers [32]. The Norwegian boiling heavy water research reactor in Halden provided heat that would have otherwise been dissipated into a nearby river to a paper mill until the reactor was closed in 2018 [28].
@@ 2.1.3 Hydrogen Production
The use of hydrogen in industrial processes, but also for energy storage, is becoming increasingly relevant in decarbonization efforts [36, 34]. In general, there exist two ways in which nuclear energy could be utilized for hydrogen production. Either via heat provision for steam methane reforming [37] and for other thermo-chemical processes such as sulfur-iodine hydrogen production at temperatures greater than 750°C [38, 39], or via the generation of electricity for an electrolyzer, which would lead to further competition with renewables and raise efficiency concerns [38, 40]. Pinsky et al. [41] assess the applicability of so-called nuclear hybrid systems for hydrogen production and conclude that most currently available technologies for which nuclear could become competitive remain commercially unavailable. This assessment is supported by the lack of commercial nuclear hydrogen production projects, such as the scrapped U.S. Department of Energy project [37, 42].
@@ 2.1.4 Other Applications
Other non-electrical nuclear power applications include marine propulsion technology, most recently applied in Russia for the Akademik Lomonosov [23], and in nuclear submarines [43]. Other nuclear propulsion devices are used in outer space [26]. Seawater desalination is a particular use case for nuclear power applied at several locations [44]. Different desalination processes require different forms of energy, such as electricity (reverse osmosis) or heat (multi-effect distillation or multi-stage flash) [45], and are thus applicable to different reactor types or configurations. Required temperatures range from 200 to 300 °C [24]. Operational experience is limited to Japan, where since 1978, one BWR and seven PWRs use district heat for small seawater desalination plants, to Kazakhstan, where the first nuclear desalination facility at Aktau operated a sodium-cooled fast reactor from 1973 to 1999, to several Russian civilian nuclear-powered ships, operated since the 1980s, and to the Diablo Canyon power plant in California [46, 32].
@@ 2.2 Nuclear Technologies and Availability
The nuclear industry has been in decline for several decades, especially in OECD countries, as only a few projects were realized, which were behind schedule, well over budget, and resulted in a decline of major industry actors (e.g., the bankruptcy of Westinghouse) [12]. Most discussed nuclear technologies are either unavailable or economically disadvantageous compared to other low-carbon technologies, and thus, new-builds are mainly realized in state-owned spheres, especially in China and Russia [4, 23]. Regardless, ambitious expansion plans for nuclear power are being publicly announced [11], and nuclear power continues to play a significant role in energy scenarios [1]. While the global nuclear fleet is dominated by high-capacity LWRs, electrical and non-electrical applications are envisioned to be implemented by using LWRs with capacity of less than 300 MWel (referred to here as "small-modular reactors" (SMRs)) and so-called "new reactor concepts" (SNR), also referred to as "advanced" or "Generation IV" reactors, that comprise several non-light water reactor concepts [7, 29]. These alternative concepts must also fulfill four prerequisites to be viable for large-scale deployment in a future decarbonized energy system. They must be 1) affordable, 2) economically competitive with other low-carbon technologies, 3) socially acceptable, and 4) available on a commercial scale [13]. In the following, we briefly discuss these prerequisites for often discussed reactor types, while omitting questions of social acceptance. See Appendix B for more technical details and information on past and ongoing research and development for the here-discussed non-light water reactor technologies.
@@ 2.2.1 Light Water Reactors
High-capacity light water reactors (LWR), i.e., reactors exceeding capacities of 300 MWel, referred to also as "Generation III(+)" reactors, dominate global reactor fleets today [23]. These reactors are usually operated as "baseload" plants with relatively high capacity factors and are mostly limited to electricity production. As discussed in Section 2.1, individual plants have also produced district or process heat. Currently, the construction of new plants is limited mostly to Russian and Chinese projects and will have to compensate for aging reactor fleets facing closure [23]. Projects in Europe and the U.S. are pained by cost and construction duration escalations [3], and a limited actor base shows an industry in decline [12]. Thus, it is highly questionable whether these reactor types will be available on a sufficient scale for energy system decarbonization in the coming decades [7]. Independent and reliable cost data is available for Western (OECD) countries [16]. At current cost levels, LWRs are not economically competitive against renewable energies in a low-carbon energy system [47, 8]. Nonetheless, assuming that aging fleets were to be replaced with new reactors, flexible operation and heat provision are technically feasible and could bring benefits to a system characterized by a high degree of renewables, such as lower system operational costs, increased reactor owner revenues, and reduced curtailment of renewables [19].
@@
SMRs are envisioned to reduce costs compared to LWRs mainly due to modularization and factory production of components, facilitated regulation, passive safety systems, and shorter construction times that would reduce capital costs, amongst others [48, 5]. Further, SMRs are proposed for niche applications, such as energy-intensive operations in remote locations [49]. In 2014, Liu and Fan [50] assumed that light water SMRs could be on a high technology readiness level of 7 to 8 and could be elevated to level 9, i.e., the operation of a functioning system, within the next ten years because of experience of similar reactor concepts in the 1950s and 1960s. However, the transferability of knowledge from early concepts should be questioned, primarily due to increased safety requirements for nuclear reactors [51]. Operation and construction experience of SMRs is limited to a small number of research and prototype reactors that have, in parallel to their high-capacity counterparts, also suffered from decade-long construction times or are still in the licensing stage and remain years away from actual operation [7, 52]. Additionally, to be economically competitive with other low-carbon technologies, SMRs would first have to overcome diseconomies of scale, and even then, cost projections by manufacturers are likely overly optimistic [5]. Nonetheless, SMRs are considered for potential non-electrical applications. For example, the NuScale VOYGR is expected to be able to provide a pre-heated feedstock of heat for refinery processes at around 300 °C [45], and it is supposed to be compatible with a high-temperature electrolyzer for hydrogen production [45, 53, 37]. The only VOYGR order, however, was canceled in late 2023 due to cost escalations [54].

### v2

# Can They Compete? Cost Competitiveness of Non-Light-Water Reactors for Heat and Power Supply in a Decarbonized European Energy System

@@ Abstract
Recent pledges to triple global nuclear capacity by 2050 suggest a "nuclear renaissance," bolstered by unconventional reactor concepts such as sodium-cooled fast reactors, high-temperature reactors, and molten salt reactors. These technologies claim to address the challenges of today’s high-capacity light-water reactors, i.e., cost overruns, delays, and social acceptance, while also offering additional non-electrical applications. However, this analysis reveals that none of these concepts currently meet the prerequisites of affordability, competitiveness, or commercial availability. Our cost analysis reveals optimistic FOAK cost assumptions of 5,623 to 9,511 USD per kW, and NOAK cost projections as low as 1,476 USD per kW. At FOAK cost, the applied energy system model for Europe in 2040 includes no nuclear power capacity, and thus indicates that significant cost reductions would be required for these technologies to contribute to energy system decarbonization. In lower-cost scenarios, reactors capable of producing high and medium temperature heat compete with other technologies and dominate the system once costs fall below 5,000 USD per kW. Electricity shares reach current levels of $\approx 20\%$ once costs are reduced to 3,000 USD per kW or less We conclude that, for reactor capacities to increase significantly, a focus on certain technology lines and streamlined regulation in necessary. Further remaining technological challenges, e.g., new waste streams, must be resolved.
@@
Analysis of historical and future non-electrical applications of nuclear power
Detailed cost analysis for the proposed non-light water reactor concepts
Energy system modeling of nuclear cogeneration competing with renewables
@@ 1 Introduction
There is a broad consensus that decarbonizing energy systems is necessary to limit global warming and reduce the potentially hazardous effects of climate change [1]. However, extensive debates exist on the means to this end, especially regarding the technology mix for a cost-efficient and executable solution. In these debates, nuclear power is often considered as one of the low-carbon technologies that will play a significant part in global energy system decarbonization [2, 1, 3], albeit existing uncertainties regarding the cost-efficiency and availability of this technology [4, 5, 6, 7, 8, 9]. In the last decade, these prospects have launched a narrative of a "nuclear renaissance" coinciding with the promotion of so-called novel reactor concepts, or "advanced" reactor technologies, mostly based on non-light water designs [10, 11]. This is supported by recent governmental proclamations to triple the globally installed nuclear power plant capacity by 2050 [12].
@@
These ambitious expansion targets stand against several challenges that the industry is currently facing [8], including cost escalations when building new high-capacity light water reactors (LWR) [4, 5], uncertainty regarding the availability of so-called "small modular reactors" (SMR) and other "advanced" reactor concepts [6] as well as a limited actor base [13]. These challenges raise concerns about whether the nuclear industry will be able to live up to the high expectations raised by above-mentioned announcements.
@@
Currently, several "novel" concepts are being developed. According to a 2023 report by the U.S. "Committee on Laying the Foundations for New and Advanced Nuclear Reactors" [14, p.43], these concepts must fulfill four conditions to be competitive in an energy system with a high share of fluctuating renewables: They must be 1) "affordable for owner investment without ’betting the company’", 2) "economically competitive with other technologies […]", 3) "socially acceptable", and 4) "commercially available." This coincides with required overnight construction costs in the range of 2,000 to 6,000 USD per kW and the establishment and diffusion of non-electrical uses for nuclear power plants [15, 16, 14]. However, there is a considerable discrepancy between cost assumptions or projections and actual costs of high capacity LWR projects, especially in OECD countries [9]. For SMR and non-LWR concepts, limited operational experience hinders the availability of accurate cost and other performance-related data [8]. Especially for SMR concepts, manufacturers’ cost projections appear somewhat optimistic compared to conventional production theory [6, 17]. Nonetheless, smaller sizes, non-electrical uses, and other benefits of nuclear, such as reliability and flexibility, could increase the competitiveness of both SMR and other concepts [18, 19, 20, 21].
@@
Previous studies on the potential viability of nuclear power in decarbonized energy systems limit their assessments to existing nuclear technologies (i.e., high capacity LWRs) and electricity production [22, 23, 9], consider only a limited number of low-carbon flexibility options [2, 24, 15, 25], limit their analysis to cost parameters [17, 6, 26] or focus on a single country considered as an energy island [24, 2, 27, 28, 29]. Some literature does refer to "advanced" reactor technologies with links to (thermal) storage options (e.g., [2, 30]). However, to our knowledge, a gap exists regarding the detailed assessment of the potential of non-electrical applications for nuclear power in an integrated energy system. Further, a model-based consideration of the potential non-electrical applications of different nuclear reactor concepts is missing due to the limitation of previous studies on a single respective reactor design.
@@
Thus, in this work, we assess the potential of SMR and non-LWR concepts in future energy systems, including non-electrical applications, by following a three-step approach. First, based on an extensive literature assessment, we investigate potential non-electrical applications of nuclear power, e.g., for heat applications, and discuss technology availability. Second, we discuss the potential costs for each technology based on an analysis of 65 references. Third, based on the prior steps, we investigate the economic efficiency of these technologies by integrating them into a European multi-vector energy system model [31] with the ability to generate both electricity and heat. Assuming non-LWR technologies become commercially available within the next decede, we find that costs must be reduced substantially, i.e., well below current LWR project costs in Europe. These reactor concepts are applied mainly for heat provision for which competitive low-carbon technologies are similarly expensive.
@@
The remainder of this work is structured as follows. Section 2 provides the main results of the literature assessment on use cases and technology availability, and of the cost data analysis. Section 3 introduces the applied energy system model and relevant assumptions and data. The results of the energy system model are presented in Section 4, and are discussed in Section 5. Section 6 concludes. Further background information is provided in the Appendix.
2 Background on Nuclear Reactor Technologies
@@
This section gives an overview of potential non-electrical use cases for nuclear power and available reactor concepts. Further, we provide an assessment of cost projections for these currently commercially unavailable reactor concepts.
@@ 2.1 Potential Non-electrical Applications of Nuclear Power
Despite the dominance of LWRs in today’s global nuclear fleet, which is mainly used to generate electricity [32], the cogeneration of electricity and heat has been the subject of literature and technology development for decades, as in general, nuclear reactors are already "heat generation devices" [33, p. 23]. Csik and Kupitz [33] further differentiate between reactors intended for heat generation only and those that can also produce electricity, allowing some degree of flexibility. For example, in a recent document, the European Commission stresses the baseload and flexible nature of nuclear power plants that provide "heat to end users in the industry and household sectors […] today" [3][p.22]. However, cogeneration experience remains limited to individual applications; see Appendix A. This results from the low-level output temperature from LWRs, which limits economic use cases. However, the output temperature could theoretically be raised at the expense of electricity generation [34].
@@
Figure 1: Potential industrial applications for existing and future reactor types, own depiction based on [36, p. 12]
@@ 2.1.1 District Heat
Nuclear power reactors have been supplying low-temperature heat up to 100°C for district heating for decades [39, 40]. While there are issues related to acceptance, e.g., addressing the proximity to populated areas [14], regulation, e.g., regarding the establishment of emergency planning zones [40, 20], as well as technology, e.g., the challenge of efficiency losses when transferring heat over long distances [34, 19], SMRs and non-LWR concepts are thought to be especially applicable to this use case [40, 29]. There have been several attempts to design reactors specifically for district heating in Sweden, Finland, and China [39, 40]. Mostly in Eastern Europe, there exists decade-long experience with district heating via nuclear cogeneration, for example, at the Kozloduy plant in Bulgaria, the Bohunice V-2 plant in Slovakia, or the Paks plant in Hungary. The heat is used for small demand centers nearby [41, 38]. Additionally, in 2019, two Westinghouse AP1000 reactors began providing district heat to the city of Haiyang in China [42].
@@ 2.1.2 Process Heat for Industry
Decarbonizing the provision of heat for industrial processes such as cement production and steel manufacturing is challenging as most processes require high temperature levels of several hundred °C and a near-constant provision of energy [43]. For decades, it has been argued that nuclear power can provide both the required heat and capacity. For example, current high-capacity LWRs can theoretically produce heat ranging from 200 to 300°C, applicable for pulp, paper, and textiles manufacturing or seawater desalination (see Section 2.1.4), and advanced gas-cooled reactors currently operating in the U.K., can provide heat up to 650°C, a temperature level relevant for the chemical industry [33]. So far, however, experience is limited to individual, mostly low-capacity demonstration projects that use mainly excess heat from electricity generation. Examples include the Calder Hall gas-cooled reactors, built in the 1950s in the U.K., which provided steam to the nearby Windscale reprocessing plant, or the Gösgen plant in Switzerland, which has been supplying heat to a nearby cardboard factory using 1% of its steam since 1979 [44]. In Canada, the Bruce heavy-water nuclear plant provided heat for its own usage but also to a greenhouse, an ethanol plant, and a plastic film production plant, amongst others, covering distances of over six kilometers [41]. The Norwegian boiling heavy water research reactor in Halden provided heat that would have otherwise been dissipated into a nearby river to a paper mill until the reactor was closed in 2018 [37].
@@ 2.1.3 Hydrogen Production
The use of hydrogen in industrial processes, but also for energy storage, is becoming increasingly relevant in decarbonization efforts [45, 43]. In general, there exist two ways in which nuclear energy could be utilized for hydrogen production. Either via heat provision for steam methane reforming [46] and for other thermo-chemical processes such as sulfur-iodine hydrogen production at temperatures greater than 750°C [47, 48], or via the generation of electricity for an electrolyzer, which would lead to further competition with renewables and raise efficiency concerns [47, 49].
Pinsky et al. [50] assess the applicability of so-called nuclear hybrid systems for hydrogen production and conclude that most technically feasible approaches for which nuclear could become competitive remain commercially unavailable. This assessment is supported by the lack of commercial nuclear hydrogen production projects, such as the scrapped U.S. Department of Energy project [46, 51].
@@ 2.1.4 Other Applications
Other non-electrical nuclear power applications include marine propulsion technology, most recently applied in Russia for the Akademik Lomonosov [32], and in nuclear submarines [52]. Other nuclear propulsion devices are used in outer space [35]. Seawater desalination is a particular use case for nuclear power applied at several locations [53]. Different desalination processes require different forms of energy, such as electricity (reverse osmosis) or heat (multi-effect distillation or multi-stage flash) [54], and are thus applicable to different reactor types. Required temperatures range from 200 to 300 °C [33]. Operational experience is limited to Japan, where since 1978, one BWR and seven PWRs use district heat for small seawater desalination plants, to Kazakhstan, where the first nuclear desalination facility at Aktau operated a sodium-cooled fast reactor from 1973 to 1999, to several Russian civilian nuclear-powered ships, operated since the 1980s, and to the Diablo Canyon power plant in California [55, 41].
@@ 2.2 Nuclear Technologies and Availability
The nuclear industry has been in decline for several decades, especially in OECD countries, as only a few projects were realized, which were behind schedule, well over budget, and resulted in a decline of major industry actors (e.g., the bankruptcy of Westinghouse) [13]. Most discussed nuclear technologies are either unavailable or economically disadvantageous compared to other low-carbon technologies, and thus, new-builds are mainly realized in state-owned spheres, especially in China and Russia [5, 32]. Regardless, ambitious expansion plans for nuclear power are being publicly announced [12], and nuclear power continues to play a major role in energy scenarios [1]. While the global nuclear fleet is dominated by high-capacity LWRs, electrical and non-electrical applications are envisioned to be implemented by using LWRs with capacity of less than 300 MWel (referred to here as "small-modular reactors" (SMRs)) and so-called "advanced" or "Generation IV" reactors, that comprise several non-LWR concepts [8, 38]. These alternative concepts must also fulfill four prerequisites to be viable for large-scale deployment in a future decarbonized energy system. They must be 1) affordable, 2) economically competitive with other low-carbon technologies, 3) socially acceptable, and 4) available on a commercial scale [14]. In the following, we briefly discuss these prerequisites for often discussed reactor types, while omitting questions of social acceptance. See Appendix B for more technical details and information on past and ongoing research and development for the here-discussed non-LWR technologies.
@@ 2.2.1 Light Water Reactors
High-capacity light water reactors (LWR), i.e., reactors exceeding capacities of 300 MWel, referred to also as "Generation III(+)" reactors, dominate global reactor fleets today [32]. These reactors are usually operated as "baseload" plants with relatively high capacity factors and are mostly limited to electricity production. As discussed in Section 2.1, individual plants have also produced district or process heat. Currently, the construction of new plants is limited mostly to Russian and Chinese projects and will have to compensate for aging reactor fleets facing closure [32]. Projects in Europe and the U.S. are pained by escalating costs and construction times [4], and a limited actor base shows an industry in decline [13]. Thus, it is highly questionable whether these reactor types will be available on a sufficient scale for energy system decarbonization in the coming decades [8]. Independent and reliable cost data is available for Western (OECD) countries [18]. At current cost levels, LWRs are not economically competitive against renewable energies in a low-carbon energy system [56, 9]. Nonetheless, assuming that aging fleets were to be replaced with new reactors, flexible operation and heat provision are technically feasible and could bring benefits to a system characterized by a high degree of renewables, such as lower system operational costs, increased reactor owner revenues, and reduced curtailment of renewables [22].
@@
SMRs are envisioned to reduce costs compared to LWRs mainly due to modularization and factory production of components, facilitated regulation, passive safety systems, and shorter construction times that would reduce capital costs [57, 6, 17]. Further, SMRs are proposed for niche applications, such as energy-intensive operations in remote locations [58, 59, 60] or microgrids [61, 62]. In 2014, Liu and Fan [63] assumed that light water SMRs could be on a high technology readiness level of 7 to 8 and could be elevated to level 9, i.e., the operation of a functioning system, within the next ten years because of experience of similar reactor concepts in the 1950s and 1960s. However, the transferability of knowledge from early concepts should be questioned, primarily due to increased safety requirements for nuclear reactors [64]. Operation and construction experience of SMRs is limited to a small number of research and prototype reactors that have, like their high-capacity counterparts, also suffered from decade-long construction times or are still in the licensing stage and remain years away from actual operation [8, 65]. Additionally, to be economically competitive with other low-carbon technologies, SMRs would first have to overcome diseconomies of scale, and even then, cost projections by manufacturers are likely overly optimistic [6]. Nonetheless, SMRs are considered for potential non-electrical applications. For example, the NuScale VOYGR is expected to be able to provide a pre-heated feedstock of heat for refinery processes at around 300 °C [54], and it is supposed to be compatible with high-temperature electrolyzers for hydrogen production [54, 66, 46]. The only VOYGR order, however, was canceled in late 2023 due to cost escalations [67]. Other projects, such as the Argentinian CAREM-14 reactor, have been scrapped, or are facing increasing delays, like the Ontario Power Generation’s BWRX-300 project in Canada [68].